find the roots using bisector method Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

2x^5+x^4-2x-1=0
1
Expert's answer
2020-10-27T19:08:40-0400
"f(x)=2x^5+x^4-2x-1"

"a=0, b=1.5"

"f(a)=f(0)=2(0)^5+(0)^4-2(0)-1=-1"

"f(b)=f(1.5)=2(1.5)^5+(1.5)^4-2(1.5)-1=16.25"

"f(a)f(b)=f(0)f(2)=-1(16.25)<0"

"x_n=\dfrac{a_n+b_n}{2}"

"a_{n+1}=x_n, b_{n+1}=b_n, f(a_n)f(x_n)\geq0"

"a_{n+1}=a_n, b_{n+1}=x_n, f(b_n)f(x_n)\geq0"

"|f(x_n)|\leq \varepsilon=>answer =x_n"

"\begin{matrix}
 n & x_n & f(x_n)\\
 0 & 0.75 & -1.708984375 \\
 1 & 1.125 & 1.955871582 \\
 2 & 0.9375 & -0.654130936 \\
 3 & 1.03125 & 0.401133597 \\
 4 & 0.984375 & -0.181243243 \\
 5 & 1.0078125 & 0.095348399 \\
 6 & 0.99609375 & -0.046479699 \\
 7 & 1.001953125 & 0.023536861 \\
 8 & 0.9990234375 & -0.011693977 \\
 9 & 1.00048828125 & 0.005865577 \\
 10 & 0.999755859375 & -0.002928138 \\
 11 & 1.0001220703125 & 0.001465231 \\
 12 & 0.99993896484375 & -0.000732325 \\
 13 & 1.000030517578125 & 0.000366235 \\
14 & 0.9999847412109375 & -0.000183099 \\
15 & 1.00000762939453125 & 0.000091554 \\
\end{matrix}"


"\big|\dfrac{0.999755859375-1.00048828125}{1.00048828125}\big|\cdot100\%\approx"

"\approx0.073\%>0.05\%"

"\big|\dfrac{1.0001220703125-0.999755859375}{0.999755859375}\big|\cdot100\%\approx"

"\approx0.037\%<0.05\%"




"\big|\dfrac{1.000030517578125-0.99993896484375}{0.99993896484375}\big|\cdot100\%\approx"


"\approx0.009\%<0.05\%"

Root is "1.000122"